# Text to model via SysML: Automated generation of dynamical system computational models from unstructured natural language text via enhanced System Modeling Language diagrams

### v1

@@ 8 Conclusions
The key enabler of the proposed approach has been the enhancement of the automated SysML diagrams initially proposed in [6] by introducing: (i) a spelling correction preprocessing step to improve component name recognition; (ii) coreference resolution to ensure consistent mapping of entities within BDD diagrams; and (iii) an automatic extraction of component attributes strategy for capturing essential information for modeling each component (e.g. information on the properties, behaviors and constraints of each component in the system). Moreover, a strategy for automatically generating an object-oriented computational model whose structure reflects that of the underlying dynamical system starting from the generated SysML diagram has been proposed. The first step consists of converting the BDD diagram into code that models the components of the dynamical system. By leveraging summarization tools and context extraction, informative docstrings for classes and functions, were generated facilitating further modification and investigation by engineers. The proposed computational model generation step uses the generated code to run simulations of the dynamical system across various input conditions.

### v2

@@ 8 Conclusions
The key enabler of the proposed approach has been the enhancement of the automated SysML diagrams initially proposed in [6] by introducing: (i) a spelling correction preprocessing step to improve component name recognition; (ii) coreference resolution to ensure consistent mapping of entities within BDD diagrams; and (iii) an automatic extraction of component attributes strategy for capturing essential information for modeling each component (e.g. information on the properties, behaviors and constraints of each component in the system). Moreover, a strategy for automatically generating an object-oriented computational model whose structure reflects that of the underlying dynamical system starting from the generated SysML diagram has been proposed. The first step consists of converting the BDD diagram into code that models the components of the dynamical system. By leveraging domain and expert knowledge (in the form of equation implementation templates), summarization tools and context extraction, informative docstrings for classes and functions were generated. The proposed computational model generation step uses the generated code to run simulations of the dynamical system across various input conditions.